Personnel On-Scene: 02/05/08 EPA-1, PADEP-1, ERRS-8, Hitchcock-2, TGA-1, Other-2, Security-1 (24 hr) 02/06/08 EPA-1, START-1, ERRS-8, TGA-3, Other-2, Security-1 (24 hr) Weather: 02/05/08 Low of 51° F, high of 56° F. Overcast, 0.00 inches of precipitation, wind SW 5-9 mph. 02/06/08 Low of 39° F, high of 43° F. Overcast, 0.01 inches of precipitation, NNE 3-15 mph.
02/05/08 EPA’s OSC Dennis Matlock mobilized to site for oversight of activities and site entry. OSC Matlock demobilized.
PADEP Mariruth Hoffman mobilized to site and assisted the EPA in oversight of activities and site entry and exit. PADEP Hoffman also changed out filters in the High Volume Air Samplers. PADEP Hoffman demobilized.
PADEP Hoffman escorted Michael Medina of Hitchcock Industries, Inc and his contractor on site to identify their material.
ERRS contractor Guardian Environmental, Inc held its morning safety meeting and discussed the days planned activities, areas of concern and safety protocol. ERRS continued to react material from Building 2 in the reaction pit. ERRS added the contents of drums marked as residual waste into the pit were water was added and the material was allowed to react. ERRS also had a contractor on site, T.G. Adams and Associates. The RAD drums will be screened and labeled prior to being shipped off site for disposal.
START was off site due to mandatory training.
02/06/08 EPA’s OSC Jack Downie mobilized to site for oversight of activities and site entry and exit.
ERRS contractor Guardian Environmental, Inc held its morning safety meeting and discussed the days planned activities, areas of concern and safety protocol. ERRS continued reacting material from Building 2 in the reaction pit. The material continues to give off heat and gas. ERRS personal are using full face respirators any time they are either working the pit or in the vicinity of the reaction pit.
START mobilized to site and assisted the OSC with written and photographic documentation.
|